# Multiple Streams of Relation Extraction: Enriching and Recalling in Transformers

### v1

@@ 4.1 Which positions are sufficient for relation extraction?
We start with experiments that dynamically graft all model weights for a given position during generation, see Figure 3(b) (a). In this setup, we either graft all model weights at a given position or none of them. Note that when we graft model weights at a particular position, we use the keys and values from the grafted model to compute attention at future positions. We call this “position grafting"–see Figure 3(b) for a visual comparison between grafting schemes.
Our position grafting results show that information flows from entities to the final token, confirming results from Geva et al. (2023). We present results for top-5 accuracy on relation extraction on all four tested models in Figure 2. See §4.1.2 for why we use top-5 accuracy.
We find that grafting only the first-entity and the last token from the finetuned to the pretrained model essentially recovers top-5 performance for all four models tested. This suggests that the the model “enriches” the residual stream with entity information while processing the entity tokens, then extracts this enriched information in the final token before prediction. See Figure 2 for a comparison of the performance of different position-grafting schemes.
Surprisingly, we also see that, in some cases, patching only the first entity or only the last token is sufficient to recover good relation extraction performance. This implies two things: (1) If an entity is “enriched" with relation information, generic mechanisms can extract that relation and (2) “recall" mechanisms can extract relation information from unenriched entities.
@@
While the “recall" and “enrichment" pathways individually have worse top-5 accuracy than when combined, for several models and sentence templates, a single pathway can be sufficient for relation extraction for some examples. In Gemma-1.1, the “recall" pathway alone achieves 53% top-5 accuracy on relation extraction (compared to a finetuned baseline of 100%) and the “enrichment" pathway for GPT2-XL reaches 28% top-5 accuracy. We conduct the same experiments on the Fake Movies, Fake Actors dataset and the Real Movies, Real Actors (Shuffled) dataset as well and find similar results (see Appendix C).
Top-5 Accuracy on Test Headline for Position-Level Patching
@@
4.1.1 Necessity for Relation Extraction
While we can show that grafting at the first entity and the final token position are sufficient to recover finetuned model performance, that does not rule out other pathways for models to extract relation information. To test this, we graft the complement of the first entity and the last token (i.e., all positions except the first entity tokens and the final token)—this results in near-zero top-k accuracy performance for all models, comparable to that of the pretrained model; see Figure 2.
We also graft the complement of the last token and the complement of the entity tokens. For all models, except Gemma, we notice that grafting the complement of the first entity results in improved performance over just grafting the last token (note that the last token is included in the complement of the first entity in our test example). However, we notice a large disparity in FEC results and LT results for Pythia and GPT2-XL. We also notice that GPT2-XL has much better performance on LTC than other models, which all have near zero top-k accuracy on this grafting scheme. See the discussion in §5 for more.
4.1.2 A Note on Top-5 Accuracy
In Figure 2, we examine the top-5 accuracy for the correct relationship token (we score the example as correct if the desired relationship token is in the top 5 choices of the next token sampling distribution). We choose top-5 accuracy since models are sometimes uncertain about which part of speech to output (i.e. sometimes they want to output the movie title instead of an actor name, or they want to output high probability tokens like “ the"). Additionally, models sometimes have multiple plausible tokenizations of an actor’s name in the top 5 (e.g. " R", " Rob", "Rob", " Robert"). This is a result of the open-endedness of our setup (simply training the model on new relations and then attempting to query the knowledge in grafted models). If a model has the correct token in the top 5 choices, we consider the model to have extracted the relation correctly. We also hypothesize that a mismatch between the pretrained and SFT model features sometimes causes predictions to regress back to the prior (token frequency unconditional on the prompt), which can result in the model defaulting to high frequency tokens. We provide additional results for token rank, which can give more fine-grained understanding, in Section C.1.2. Investigating the mechanism by which models default to the unconditional prior is a promising direction for future work.
@@ 4.2 Is it the position or the token that matters?
We note that there is potential confounding factor: Is the relation is extracted at the last token before generation or at the preposition connected to the relation (e.g., stars in)? In the test headline (see Table 1) these are always the same token. We hypothesize that the relation extraction may occur on relation tokens (e.g. “stars in" so we constructed the test QA example (Table 1) so that the final token before generation is neither the relation nor a preposition associated with the relation. We see similar results in the two test sentences: the last token is responsible for relation extraction, even when it is neither a preposition nor a relation. We present only the results for Llama3 for brevity in Figure 3(a) and present additional results for other models in Section C.1.1
Figure 3: (a) Results for Llama3 on the test QA example with a verb as the final token showing similar results to the test headline where a preposition is the final token. (b) A schematic showing the different dynamic weight grafting schemes used in our experiments. Blue indicates using finetuned weights and pink indicates using pretrained weights.)
4.3 Can we localize “recall" relation extraction to model components?
@@
Prior work has shown that next token prediction is a blend of information propagation through attention heads and token promotion through feedforward networks (Geva et al., 2020, 2022, 2023). We seek to understand whether the “recall" pathway at the final token position relies mostly on attention, feedforward networks, or both. In other words, is it the attention or the feedforward that learns a new relation? (Of course, it can be both.)
Training on disjoint data to localize relation extraction
We attempt to localize relation extraction by only grafting specific model components between models trained on the task and models trained on the actual relation. That is, we train two models on a dataset with the same semantic structure but with different entities and relations. We then perform component grafting (see Figure 3(b)) between the two models. This way, we can see which components are responsible for general task functionality and which components are responsible for relation extraction.
@@
We start by leveraging the reversal curseBerglund et al. (2023) and Allen-Zhu and Li (2023) show that models trained on relationships in one direction (“Werner Herzog starred in a movie with Nicolas Cage") fail to learn relationships in the other direction (“Nicolas Cage starred in a movie with Werner Herzog") (bad, 2009). We exploit this effect to study relation extraction by grafting weights from a model trained on both directions of a relationship (e.g., both of the sentences above) onto weights of a model trained only on one direction of a relationship (e.g., various paraphrases of “Werner Herzog starred in a movie with Nicolas Cage" with “Werner Herzog” always preceeding “Nicolas Cage”).
Recall that a Transformer block is described by:
$\text{Block}(x)=\text{LN}\bigg{(}x+\text{ATTN}\big{(}\text{LN}(x)\big{)}\,O+\text{FFN}\Big{(}\text{LN}\big{(}x+\text{ATTN}\big{(}\text{LN}(x)\big{)}\,O\big{)}\Big{)}\bigg{)}$ (2)
where LN is the Layer Norm, ATTN is the attention, FFN is the feedforward, and $O$ is the output matrix for multi-headed self-attention. Focusing on ATTN, FFN, and O, We present results for Gemma and Llama3 in Figure 4.333GPT2-XL and Pythia had much weaker “recall” results than Gemma and Llama—we see this in the reversal setting as well, so we present results for GPT2-XL in Section C.6.
@@
Figure 4: We graft weights from models finetuned on both directions of a symmetric relationship (actors starring in a movie together) at the last token to see which model components are responsible for relation extraction. For models with an effective “recall" pathway (Gemma & Llama), we see that the O matrix and the feedforward networks in the last quarter of the model recover most of the performance.
In Figure 4, we see that grafting the O matrix and the full FFN nearly recovers the results of grafting the full attention mechanism and the full FFN. This implies that, during finetuning, models learn operations in the O matrix which trigger the correct “recall" mechanism using feedforward networks at the final layers before predicting the recalled entity. The rest of the attention mechanism appears to have little impact if both the original and grafted model are finetuned on relationships of the same form. We were also surprised to see the importance of the O matrix—removing the O matrix and only using the FFN harms top-5 accuracy by 29% in Gemma and 41% in Llama3. We also note that grafting the O+FFN performs better than grafting the full layer with Gemma. We again hypothesize that this is related to shifts in model uncertainty—we notice that, when the model is uncertain (i.e. assigns low probability to the most likely next token) different patching schemes will impact the part of speech that the model weights highly (i.e. a name vs. a movie title vs. punctuation, etc.). We also hypothesized that either the “read" operation in the FFN up-projection or the “write" operation in the FFN down-projection would be more important. Instead, we see that both recover some relation extraction performance when paired with the O matrix, but both are necessary for good relation extraction recovery.
Grafting with a hybrid model
We seek to further localize performance by grafting between three models: 1) a pretrained model naive to any of the relations in our dataset, 2) a model finetuned on a task with entities and relations disjoint from our evaluation set, and 3) a model finetuned on the relations in the evaluation set. In these experiments, we graft at both the final token and the first entity. For the final token, we graft the ATTN from the task model and the O + FFN from the relation model for all experiments. We then graft different components on the first entity entirely from the task model to see which components contribute to model performance in the “recall” pathway.
Top-5 Accuracy on Test Headline for Component-Level Grafting
Figure 5: We graft weights from both a task model and a relation model onto a pretrained model, which we refer to as a hybrid model. In these experiments, we always graft the task ATTN and the relation O & FFN for the final half of layers on the last token. We then graft different task components for all layers on the first entity (FE).
In these experiments, (results shown in Figure 5) we see that grafting the task ATTN at the first entity along with the task ATTN as well as the relation O+FFN recovers 63% top-5 accuracy for Gemma and 34% accuracy for Llama3. Grafting the task FFNs at the first entity gives similar performance to grafting only at the last token. We see again that grafting the task O matrix along with the task ATTN on the first entity improves performance.
@@ 5 Discussion
We use dynamic weight grafting to show that, when models undergo supervised finetuning on new relation information, the entity tokens and the final token position are where relation extraction occurs–either alone can be sufficient, but at least one of them is necessary to extract relations.
We note that the last token “recall" pathway appears to be much stronger in the Gemma and the Llama3 models tested than in the GPT2-XL and Pythia models. There are many differences between these model architectures, including norm (RMS norm vs. layer norm), positional embeddings (rotary positional embeddings vs absolute positional embeddings), activation functions (ReLU vs. GeGLU vs. SwiGLU), embeddings (tied vs. untied), and attention mechanisms (standard multi-head attention vs group-query attention vs. multi-query attention). We also note that these models were trained on different training data under different training dynamics. See Table 2 for a more detailed comparison between models. We hypothesize that the more recent models have more expressive attention mechanisms that allow for better independent extraction of relation information, even on unenriched entities.
Since we are finetuning small models on synthetic data, we saw issues with catastrophic forgetting, so we trained models with less aggressive learning rate and removed weight decay (Lubana et al., 2022) while also including supplemental training examples from openwebtext and IMDB movie reviews. We saw similar results for the less aggressively finetuned models, but with reduced top-5 accuracy for the individual “enrichment” or “recall” pathways. See Section C.4 for more details.
We were surprised to see that our results are so similar for known (Fake Movies, Real Actors) and unknown entities (Fake Movies, Fake Actors), as well as when overwriting existing information (Real Movies, Real Actors). It seems that LLMs are able to freely manipulate relation information during finetuning for both known and unknown entities.
We also note that our reversal curse experiment discussed in 4.3 show slightly different results than Geva et al. (2023) on the role of attention and feedforward networks in the extraction of relation information. Geva et al. (2023) show that knocking out attention is more harmful to relation extraction than knocking out feedforward networks. Our results show that, in a setting where a model has already learned how to do a specific relation extraction task, the O matrices and the feedforward networks at the final token position are nearly sufficient to recover relation extraction performance as long as the model has task-specific attention functionality. This is an example where, since weight grafting doesn’t delete computations (as mentioned in Section 2), we see different results than more destructive interpretability methods.
@@
We leave the alternative entity enrichment pathway present in Pythia and GPT2-XL–see 4.1.1–unexplored in this work. We hypothesize that other token positions beyond the last are also able to extract relation information that is then extracted at the final token position. Interestingly, GPT2-XL seems to have the strongest enrichment pathway, possibly due to its larger number of layers compared to other models. Additionally, while we localized relation extraction to specific model components in some settings, we did not attempt to interpret those components. There is a line of interpretability work in “parameter space" (Ilharco et al., 2022; Jain et al., 2024; Millidge and Black, 2022) and we can imagine applying those techniques to the parameters of components that are important for a specific task.
@@
Our work focuses on a synthetic relation extraction task, potentially limiting its scope and generalization to other relation extraction settings with more complex sentences or more varied finetuning data. Additionally, we operationalize the “success" of relation extraction using top-k accuracy or the token rank of the correct relation during next token prediction–it’s possible that models “know" information in a way that doesn’t impact next token prediction, and our methods do not account for this. Additionally, there is a combinatorial explosion of possible grafting schemes and our experiments only explore a subset of possible grafts. While we try to rule out several failure modes for other methods of relation extraction, it’s possible that model features interact and “cancel" in surprising ways, so there may be other hidden ways of extracting relations. Our dynamic weight grafting procedure sometimes causes models to allocate most of their probability mass to common tokens like “ the" or punctuation–we hypothesize that this is because the blend of features created during weight grafting may be making the model uncertain about the next token so it reverts to a prior on common tokens.
Broader Impact
LLMs are capable of revealing sensitive and personal information (Pearson, 2025). Our work focuses on mechanistically understanding relation extraction, which could then be used to undo safeguards on existing open source models in order to extract personal information that models were trained on (but do not currently output due to safety-tuning).

### v2

@@ 3.1 Which positions are sufficient for relation completion?
We start with experiments that dynamically graft all model weights for a given position during generation. We call this “position grafting" (see Figure˜1(b) for a visual comparison between grafting schemes). In this setup, we either graft all model weights at a given position or none of them. Note that when we graft at a particular position, we use the keys and values from the grafted model to compute attention at that position, even if future positions use pretrained model parameters. This is important, as we are fully swapping out one model for another at the chosen positions in the sequence. Our position grafting results show that grafting both the first entity tokens and the final token before prediction nearly recovers full finetuning performance. We present results for top-5 accuracy on relation completion on all tested models in Figure˜2. See §3.1 for why we use top-5 accuracy.
In some cases, we also see that grafting only the first entity tokens is sufficient to recover some top-5 performance (similar to previous results showing that models enrich entity representations early in the sequence and carry that information forward (Meng et al., 2022a; Geva et al., 2023; 2022)). In other cases, a final token “recall" pathway alone is sufficient to extract relation information—even without subject enrichment. This indicates that the later layers of a finetuned model contain a mechanism to recall information from finetuning, even in response to a representation that does not contain information from the finetuning set.
@@
While the “recall" and “enrichment" pathways individually have worse top-5 accuracy than when combined, for several models and sentence templates, a single pathway can be sufficient for good relation completion performance. In Gemma-1.1, the “recall" pathway alone achieves 53
A Note on Top-5 Accuracy
In Figure 2, we examine the top-5 accuracy for the correct relationship token (we score the example as correct if the desired relationship token is in the top five choices of the next token sampling distribution). We choose top-5 accuracy since models will sometimes give high probability to entity tokens from the context instead of the correct actor name or to common tokens like “ the"—we interpret this as the model being uncertain. Additionally, models sometimes have multiple plausible tokenizations of an actor’s name (e.g. " R", " Rob", "Rob", " Robert"). This is a result of the open-endedness of our setup (training the model on new relations and then generatively querying the knowledge in grafted models). We note that our results are a lower bound on the ability of the model to retrieve correct relation information (Jiang et al., 2020), and that the choice of $k$ for top $k$ accuracy changes the scale of the accuracy results, but not the overall pattern (see Section˜C.2 for results with additional choices of $k$).
If a model has the correct token in the top five choices, we consider the model to have correctly retrieved the relation information. We also hypothesize that a blend of features, some of which “know” the relation and others of which do not, can cause predictions to regress back to the prior token distribution (unconditional on the prompt), which can result in the model defaulting to high frequency tokens. To give a more fine-grained understanding, we provide additional results for token rank in Section˜C.1.2. Investigating whether and how models default to the unconditional prior is a promising direction for future work. We also test additional values for k. We see the same behavioral pattern for different values of k, but the strength of the individual “enrichment” and “recall” pathways is weaker for lower values of k. See Section˜C.2 for more details.
Top-5 Accuracy on Test Headline for Position-Level Grafting
@@ A Note on Top-5 Accuracy
3.1.1 Necessity for Relation Completion
While we can show that grafting at the first entity and the final token position are sufficient to recover finetuned model performance, this does not rule out other pathways for models to extract relation information. To test this, we graft the complement of the first entity and the last token: (FE+LT)C (i.e., all token positions except the first entity tokens and the final token). This results in near-zero top-k accuracy performance for all models, comparable to that of the pretrained model (Figure˜2).
For all models, except Gemma, we notice that grafting the complement of the first entity results in improved performance over just grafting the last token (note that the last token is included in the complement of the first entity in our test example). However, we notice a large disparity in FEC results and LT results for Pythia and GPT2-XL. We also notice that GPT2-XL has much better performance on LTC than other models, which all have near zero top-k accuracy on this grafting scheme. We also run experiments with the movie title included in the test sentence (see Section˜C.5); we find that the movie title alone is not sufficient to recover finetuned performance, but the movie title and the last token together give improved performance over the last token alone. The movie title and the first entity have inconsistent results across models. See the discussion in §4 for more.
@@ 3.2 Is it the position or the token that matters?
There is ambiguity in the setup described: Is the relation extracted at the last token before relation generation or at the preposition connected to the relation (e.g., stars in)? In the test headline (see Table˜1), these are always the same token. We hypothesize that the relation completion may occur at relation tokens so we constructed the QA test (Table˜1) so that the final token before generation is neither the relation nor a preposition associated with the relation. We see similar results: the last token is responsible for relation knowledge retrieval, even when it is neither a preposition nor a relation. We present results for these experiments in Section˜C.1.1
Figure 3: We localize the “recall” pathway of finetuned knowledge retrieval to specific model components. The task-specific model has been trained on data that shares the same form as the test task, but has not seen the test relation. The relation-specific model has been trained on the test relations. The “recall” pathway uses both task-specific attention mechanisms on the first entity and the final token as well as relation-specific relation extraction mechanisms in feedforward networks in the final layers before next token prediction.
3.3 Can we localize the “recall" pathway to individual model components?
@@
Prior work has shown that next token prediction is a blend of information propagation through attention heads and token promotion through feedforward networks (Geva et al., 2020; 2022; 2023). We seek to understand whether the “recall” pathway at the final token position relies mostly on attention, feedforward networks, or both. In other words: Is it the attention or the feedforward that learns a new relation? (Of course, it can be both.)
Component grafting
In component grafting, we graft individual components of the Transformer block at a given token position (Figure˜1(b)). Recall that a Transformer block is described by  [^3] There are model-specific subtleties to the attention and feedforward operations. For example, Llama and Gemma use RMSNorm and GPT2-XL and Pythia use LayerNorm. Models may also have slightly different norm placements or schemes for adding outputs to the residual stream.:
$\text{Block}(x)=x\;+\;\operatorname{ATTN}(\operatorname{NORM}(x))\,O\;+\;\operatorname{FFN}(\operatorname{NORM}(x+\operatorname{ATTN}(\operatorname{NORM}(x))\,O))$ (1)
where the NORM operation is either Layer Norm or RMSNorm, ATTN is the attention, FFN is the feedforward network, and $O$ is the output projection matrix for multi-headed self-attention. We focus on grafting individual weight matrices at a given token position; for example, we may graft the $W^{K},W^{Q}$ and $W^{V}$ matrices at the first entity token positions. At a high level, we conduct two component grafting experiments to localize model behavior for the last token “recall” pathway: 1) We localize relation knowledge retrieval to the output projection matrix and feedforward networks at the final token position, and 2) We show that models require task-specific attention on the first entity in order for the “recall” pathway to work. See Figure˜3 for a visual representation of the these results.
Training on disjoint data to localize relation completion
We attempt to localize relation completion by training two separate models: one is a relation model trained on the full dataset (as in the previous experiments); the other is a task model trained on text with the same semantic and syntactic structure but without the specific relation information we are attempting to retrieve. See Section˜A.2.4 for examples of the data used to train a task model vs. a relation model.
@@
In these experiments, we leverage the reversal curseBerglund et al. (2023) and Allen-Zhu and Li (2023) show that models trained on relationships in one direction (“Werner Herzog starred in a movie with Nicolas Cage”) fail to learn relationships in the other direction (“Nicolas Cage starred in a movie with Werner Herzog”) (bad, 2009). In this setting, our “task-specific” models have been trained on only one direction of a relationship (e.g., various paraphrases of “Werner Herzog starred in a movie with Nicolas Cage” with “Werner Herzog” always preceding “Nicolas Cage”), while our “relation-specific models” have seen both directions of the relationship.
Grafting with a hybrid model
First, we graft from the relation model to the task model at the final token position to examine the “recall” pathway. We present results for Gemma and Llama3 in Figure˜4.444GPT2-XL and Pythia had much weaker “recall” results than Gemma and Llama—we present results for GPT2-XL in Section C.9. In Figure˜4, we see that grafting the O matrix and the full FFN nearly recovers the results of grafting the full attention mechanism and the full FFN. This implies that, during finetuning, models learn operations in the O matrix which trigger the correct “recall” mechanism using FFNs at the final layers before predicting the recalled entity. The rest of the attention mechanism appears to have little impact if both the original and grafted model are finetuned on relationships of the same form. We were also surprised to see the importance of the O matrix—removing the O matrix and only using the FFN harms top-5 accuracy by 29
Next, we graft between three models: 1) a pretrained model (naive to any of the relations in our dataset), 2) a task model, and 3) a relation model. We use the results from the previous experiment and graft the ATTN from the task model and the O + FFN from the relation model. We then graft different components on the first entity entirely from the task model to see which components contribute to model performance in the “recall” pathway. In Figure˜5, we see that grafting the task ATTN at the first entity and, at the last token, the task ATTN and the relation O+FFN recovers 63
@@
Figure 4: We graft weights from models finetuned on both directions of a symmetric relationship at the last token to identify which components drive relation completion. For Gemma and Llama, the output projection matrix and feedforward networks in the last quarter of the model recover most of the finetuned performance.
Top-5 Accuracy on Test Headline for Component-Level Grafting (Hybrid)
Figure 5: We graft from both a task model and a relation model onto a pretrained model, creating a hybrid model. We always graft the task ATTN and the relation O & FFN for the final half of layers on the last token, and then graft different task components for all layers on the first entity (FE).
@@ 4 Discussion
We use dynamic weight grafting to show that:
  • When models undergo supervised finetuning on new relation information, the entity tokens and the final token position are where relation completion occurs—either pathway alone can be sufficient, and both nearly recover full finetuning performance.
  • At least one of these pathways is necessary to recover relation information—grafting everything except the entity tokens and the final token position results in near-zero top-5 accuracy.
  • Using component grafting (see Figure˜1(b)), we localize the relation retrieval in the “recall” pathway to the output projection matrix and the feedforward networks at the final token position as well as task-specific attention at the first entity and the final token.
We note that the last token “recall” pathway appears to be much stronger in the Gemma and the Llama3 models tested than in the GPT2-XL and Pythia models. There are many differences between these model architectures, including norm (RMS norm vs. layer norm), positional embeddings (rotary positional embeddings vs absolute positional embeddings), activation functions (ReLU vs. GeGLU vs. SwiGLU), embeddings (tied vs. untied), and attention mechanisms (standard multi-head attention vs. group-query attention vs. multi-query attention). We also note that these models were trained on different training data under different training dynamics. See Table˜2 for a more detailed comparison between models. We hypothesize that the more recent models have more expressive attention mechanisms that allow for better recovery of relation information.
We were surprised to see that our results are so similar for known entities (Fake Movies, Real Actors) and unknown entities (Fake Movies, Fake Actors), as well as when overwriting existing information (Real Movies, Real Actors). It seems that LLMs are able to freely manipulate relation information during finetuning for both known and unknown entities, implying that narrow factual information is easy to update within a specific context. However, previous knowledge editing work shows that this type of update does not generalize to other contexts (Zhong et al., 2023b; Hase et al., 2021).
Since weight grafting doesn’t delete computations (see Section˜2), we see different results than more destructive interpretability methods in localizing information retrieval. Our component grafting experiments discussed in 3.3 show slightly different results than Geva et al. (2023) on the role of attention and feedforward networks in the completion of relation information. Geva et al. (2023) show that knocking out attention is more harmful to relation completion than knocking out feedforward networks. Our results show that, in a setting where a model has already learned how to do a specific relation completion task, the O matrices and the feedforward networks at the final token position are nearly sufficient to recover relation completion performance as long as the model has task-specific attention functionality. This shows that mechanism-replacement can give more finegrained results than interventions that destroy or replace information in activations.
@@
We leave the alternative entity enrichment pathway present in Pythia and GPT2-XL–see 3.1.1–unexplored in this work. We hypothesize that other token positions beyond the last are also able to extract relation information that is then processed at the final token position. Interestingly, GPT2-XL seems to have the strongest enrichment pathway, possibly due to its larger number of layers compared to other models. Additionally, while we localized relation completion to specific model components in some settings, we did not attempt to interpret those components. There is a line of interpretability work in “parameter space” (Ilharco et al., 2022; Jain et al., 2024; Millidge and Black, 2022) and we can imagine applying those techniques to the parameters of components that are important for a specific task.
@@
Our work focuses on a synthetic knowledge retrieval task, potentially limiting its scope and generalization to other settings with more complex sentences or more varied finetuning data. Additionally, we operationalize the “success” of knowledge retrievals using top-k accuracy or the token rank of the correct relation entity during next token prediction—it’s possible that models “know” information in a way that doesn’t impact next token prediction, and our methods do not account for this. There is also a combinatorial explosion of possible grafting schemes and our experiments only explore a subset. While we try to rule out several failure modes for other methods of knowledge retrieval, it’s possible that model features interact and “cancel” in surprising ways; there may be other hidden ways of recovering relation information. Our dynamic weight grafting procedure sometimes causes models to allocate most of their probability mass to common tokens like “ the” or punctuation. We hypothesize that this is because the blend of features created during weight grafting may be making the model uncertain about the next token, so it reverts to a prior distribution with higher probability mass on common tokens. We also use smaller models in our experiments due to compute limitations; it is possible that larger models have different mechanisms when finetuned on new relation information.